# There Was Never a Bottleneck in Concept Bottleneck Models

### v2

@@ 1 Introduction
Most machine learning models operate by learning data representations—compressed versions of the input that retain the essential information needed to solve a given task [1]. However, these representations often encode information in ways that are not easily interpretable by humans. This lack of interpretability becomes especially problematic in sensitive domains such as healthcare [2, 3, 4], finance [5, 6], and autonomous driving [7, 8]. To address this issue, Concept Bottleneck Models (CBMs) have been proposed, which enforce representations to be defined in terms of a set of human-understandable concepts [9].
Formally, given a task $\bm{y}$ and a set of concepts $\bm{c}=\{c_{j}\}_{j=1}^{m}$, Vanilla Models (VMs) are trained with a single objective: their representations $\bm{z}$ should capture the information necessary to predict $\bm{y}$ accurately. CBMs extend this framework by adding a second objective: each concept $c_{j}$ must be recoverable from a designated component $z_{j}\in\bm{z}$. By enforcing this additional constraint, CBMs are claimed to offer: (i) improved interpretability of the representation space, and (ii) the ability to intervene on specific concepts by modifying $z_{j}$ and propagating these changes to the predictions.
However, CBMs are prone to a phenomenon known as information leakage [10, 11], where the representation $\bm{z}$ encodes input information that cannot be attributed to the predefined concepts $\bm{c}$. We refer to this additional information as nuisances $\bm{n}$. Information leakage raises two main concerns: (i) it undermines interpretability, since $z_{j}$ cannot be fully explained by its corresponding concept $c_{j}$; and (ii) it compromises the validity of interventions—modifying $z_{j}$ may alter not only the associated concept $c_{j}$, but also other unintended information encoded in $z_{j}$.
We argue that information leakage stems from a fundamental limitation in the current formulation of CBMs: the absence of an explicit Information Bottleneck (IB) [12] that actively constrains $z_{j}$ to exclude information unrelated to $c_{j}$. While the second objective in CBMs encourages each $z_{j}$ to retain $c_{j}$ in its entirety, it does not enforce that $z_{j}$ captures only information about $c_{j}$. In the worst-case scenario, $z_{j}$ could encode the entire input $\bm{x}$ and still satisfy this objective.
@@
To address this issue, we propose Minimal Concept Bottleneck Models (MCBMs), which incorporate an IB into each $z_{j}$. This ensures that $z_{j}$ not only retains all the information about its associated concept $c_{j}$, but also excludes any information unrelated to $c_{j}$, as summarized in Table 1. The name reflects that $z_{j}$ is trained to be a minimal sufficient statistic of $c_{j}$ [13, 14], in contrast to traditional CBMs where $z_{j}$ is optimized to be merely a sufficient statistic of $c_{j}$. As illustrated in Figure 1, this design yields disentangled representations that directly address the two shortcomings of CBMs: (i) it improves interpretability, since $z_{j}$ can be fully explained by its corresponding concept $c_{j}$; and (ii) it enables valid interventions—modifying $z_{j}$ affects only the associated concept $c_{j}$.
@@ 2.1 Data Generative Process
For this scenario, we consider inputs $\bm{x}\in\mathcal{X}$, targets $y\in\mathcal{Y}$, concepts $\bm{c}=\{c_{j}\}_{j=1}^{m}\in\mathcal{C}$ and nuisances $\bm{n}\in\mathcal{N}$, such that $p(\bm{x},\bm{y},\bm{c},\bm{n})=p(\bm{x}|\bm{c},\bm{n})p(\bm{y}|\bm{x})$, i.e., the inputs $\bm{x}$ are described by the concepts $\bm{c}$ and the nuisances $\bm{n}$, and the targets $\bm{y}$ are fully described by the input $\bm{x}$. The only difference between $\bm{c}$ and $\bm{n}$ is that the former are observed, while the latter are not or, in other words, labels on $\bm{c}$ are provided to us. We assume access to a training set $\{\bm{x}^{(i)},\bm{y}^{(i)},\bm{c}^{(i)}\}_{i=1}^{N}$, which defines the empirical distribution $p(\bm{x},\bm{y},\bm{c})=\sum_{i=1}^{N}\delta\left(\bm{x}-\bm{x}^{(i)}\right)\delta\left(\bm{y}-\bm{y}^{(i)}\right)\delta\left(\bm{c}-\bm{c}^{(i)}\right)$. The graphical model corresponding to this generative process is shown in Figure 2(a) for the case of two concepts.
@@ 2.2 Vanilla Models
In machine learning, the most commonly studied problem is that of predicting $\bm{y}$ from $\bm{x}$, which serves as a foundation for more specialized tasks. We refer to models trained to address this problem as Vanilla Models (VMs). These models typically operate by first extracting an intermediate representation $\bm{z}\in\mathcal{Z}$ from the input $\bm{x}$ via an encoder $p_{\theta}(\bm{z}\mid\bm{x})$. Subsequently, a prediction $\hat{y}\in\mathcal{Y}$ is produced from $\bm{z}$ using a task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$. Since $\bm{z}$ is intended to facilitate accurate prediction of $\bm{y}$, the mutual information between $\bm{z}$ and $\bm{y}$, denoted $I(Z;Y)$, should be maximized. In Appendix B.1, we formally show that:
@@
Figure 2(b) shows the graphical model of a Vanilla Model with a two-dimensional representation $\bm{z}$. Black edges represent the encoder $p_{\theta}(\bm{z}\mid\bm{x})$, while green edges indicate the task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$. The encoder is typically chosen to be deterministic, i.e., $p_{\theta}(\bm{z}\mid\bm{x})=\delta\left(\bm{z}-f_{\theta}(\bm{x})\right)$. However, for tractability reasons (see Section 2.4), we adopt a stochastic formulation where $p_{\theta}(\bm{z}\mid\bm{x})=\mathcal{N}\left(\bm{z};f_{\theta}(\bm{x}),\sigma_{x}^{2}I\right)$, as summarized in Table 2. The choice of task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$ depends on the structure of the output space $\mathcal{Y}$, also detailed in Table 2. The objective in Equation 1 corresponds to minimizing the cross-entropy loss when $\bm{y}$ is binary or multiclass, and to minimizing the mean squared error between $\bm{y}$ and $g^{y}_{\phi}(z)$ when $\bm{y}$ is continuous.
@@ 2.3 Concept Bottleneck Models
    Concepts $c_{j}\in\bm{c}$, are assumed to be binary. If a concept is originally multiclass with $k$ categories, it is converted into $k$ binary concepts (One-vs-Rest [15]).
@@
Since $\sigma$ is invertible, this setup permits defining $p(z_{j}\mid c_{j})\approx\sigma^{-1}(c_{j})$. However, this is ill-defined at the binary extremes, as $\sigma^{-1}(1)=-\sigma^{-1}(0)=\infty$. To address this, in practice, $p(z_{j}\mid c_{j}=0)$ and $p(z_{j}\mid c_{j}=1)$ are set as the 5th and 95th percentiles of the empirical distribution of $z_{j}$, respectively [9]. This workaround, however, introduces two crucial issues discussed in Section 5.
Table 2: Distributions considered in this work. $f_{\theta}$ is typically modeled by a large neural network while $g^{y}_{\phi}$, $g^{c}_{\phi}$ and $g^{z}_{\phi}$ by simpler neural networks. We consider $\bm{z}$ to be always continuous.
@@ 2.4 Minimal Concept Bottleneck Models
    For categorical concepts, $g^{z}_{\phi}(c_{j})=\lambda\cdot\mathrm{one\_hot}(c_{j})$. This mirrors the Prototypical Learning [16] approach where class-dependent prototypes $\{g^{z}_{\phi}(c_{j})\}_{j=1}^{m}$ are fixed.
@@
To optimize MCBMs, we combine the three previously introduced objectives, incorporating two key considerations. First, we approximate the expectations over $p(\bm{x},\bm{y})$ and $p(\bm{x},c_{j})$ using the empirical data distribution, replacing integrals with summations over the dataset. Second, to enable gradient-based optimization through the stochastic encoder, we apply the reparameterization trick [17]: $\mathbb{E}_{p_{\theta}(\bm{z}|\bm{x})}\left[\log q_{\phi}(\bm{y}|\bm{z})\right]\approx\sum_{i}\log q_{\phi}\left(\bm{y}|f^{\prime}_{\theta}\left(x,\epsilon^{(i)}\right)\right)$ and $E_{p_{\theta}(z_{j}|\bm{x})}\left[\log q_{\phi}(\hat{c}_{j}|z_{j})\right]\approx\sum_{i}\log q_{\phi}\left(\hat{c}_{j}\big|f^{\prime}_{\theta,j}\left(\bm{x},\epsilon^{(i)}\right)\right)$, where $f^{\prime}_{\theta}(\bm{x},\epsilon)=f_{\theta}(\bm{x})+\sigma^{2}_{x}I\epsilon$ (due to the choice of $p_{\theta}(\bm{z}|\bm{x})$ in Table 2), $\epsilon\sim\mathcal{N}\left(0,I\right)$ and $f^{\prime}_{\theta,j}\left(x,\epsilon\right)$ corresponds to the element $j$ of $f^{\prime}_{\theta}\left(\bm{x},\epsilon\right)$. Combining the considerations above yields the final objective for training MCBMs, as shown in Equation 6, where $\beta$ and $\gamma$ are hyperparameters. The first term corresponds to the objective used in Vanilla Models, the second term is introduced in CBMs, and the third is specific to MCBMs. Detailed training algorithms for the various cases listed in Table 2 are provided in Appendix C.
@@
Figure 2: Graphical models of the different systems described for two concepts and two-dimensional representations. We provide in Appendix A the analogous figure for $m$ concepts and $m$-dimensional representations. Inputs $\bm{x}$ are defined by some concepts $\{c_{j}\}_{j=1}^{m}$ and nuisances $\bm{n}$; and targets $\bm{y}$ are defined by $\bm{x}$ (gray arrows). Vanilla models obtain the representations $\{z_{j}\}_{j=1}^{m}$ from $\bm{x}$ through the encoder $\color[rgb]{0,1,1}p_{\theta}(\bm{z}|\bm{x})$ (cyan arrows) and solve the task $\hat{\bm{y}}$ sequentially through the task head $\color[rgb]{0,1,0}q_{\phi}(\hat{\bm{y}}|\bm{z})$ (green arrows). Concept Bottleneck Models make a prediction $\hat{c}_{j}$ of each concept $c_{j}$ from one representation $z_{j}$ through the concept head $\color[rgb]{0,0.49609375,1}q(\hat{c}_{j}|z_{j})$ (blue arrows). Minimal CBMs make a prediction $z_{j}$ of each representation $z_{j}$ from one concept $c_{j}$ through the representation head $\color[rgb]{1,0,0}q(z_{j}|c_{j})$ (red arrows).
@@ 3 Related Work
Information leakage occurs when the learned representation $\bm{z}$ encodes information outside the concept set $\bm{c}$, reducing both interpretability and intervenability [10, 11]. This arises when the Markovian assumption fails—i.e., when the target $\bm{y}$ is not fully determined by the concept set $\bm{c}$, or equivalently, $p(\bm{y}|\bm{c})\neq p(\bm{y}|\bm{c},\bm{x})$, which is typically the case in real-world scenarios [18]. Concept Embedding Models (CEMs) [19] were proposed to mitigate the accuracy–interpretability trade-off in CBMs. However, this trade-off is fundamentally limited by the chosen concept set, and CEMs may even be less interpretable than standard CBMs, as their more entropic representations tend to amplify information leakage. Recently, this critique has been further formalized: Parisini et al. [20] proposed information-theoretic benchmarks for assessing interpretability and Makonnen et al. [21] introduced new metrics grounded in usable mutual information [22]. Havasi et al. [18] also introduced Hard Concept Bottleneck Models (HCBMs), which predict $\bm{y}$ from binarized concept predictions $\hat{c}_{j}$ rather than from $\bm{z}$ (see Appendix D), thereby imposing an ad-hoc Information Bottleneck. Extensions such as Autoregressive CBMs and Stochastic CBMs [18, 23] incorporate dependencies between concepts.
@@
The Information Bottleneck (IB) [12] provides a principled way to balance preserving information about a factor with compressing the representation. In this framework, a representation is sufficient if it retains all the information about the variable of interest, and minimal if it contains only that information [24, 25, 26]. Computation of these quantities is intractable, so variational methods have been developed to derive tractable evidence bounds [27, 28].

### v3

@@ 1 Introduction
Most machine learning models operate by learning data representations—compressed versions of the input that retain the essential information needed to solve a given task (Bengio et al., 2013). However, these representations often encode information in ways that are not easily interpretable by humans. This lack of interpretability becomes especially problematic in sensitive domains such as healthcare (Ahmad et al., 2018; Xie et al., 2020; Jin et al., 2022), finance (Brigo et al., 2021; Liu et al., 2023), and autonomous driving (Kim and Canny, 2017; Xu et al., 2024b). To address this issue, Concept Bottleneck Models (CBMs) have been proposed, which enforce representations to be defined in terms of a set of human-understandable concepts (Koh et al., 2020).
Given an input $\bm{x}$ and target $\bm{y}$, Vanilla Models (VMs) are trained with a single goal: the representation $\bm{z}$ derived from $\bm{x}$ should encode all information needed to accurately predict $\bm{y}$. In many settings, additional side information—often referred to as concepts—is available, denoted by $\bm{c}=\{c_{j}\}_{j=1}^{m}$. Concept Bottleneck Models (CBMs) leverage this by extending VMs with a second objective: each concept $c_{j}$ must be recoverable from a designated component $z_{j}\in\bm{z}$. By enforcing this additional constraint, CBMs are purported to provide: (i) enhanced interpretability of the learned representation space, and (ii) the ability to perform targeted interventions on specific concepts by manipulating $z_{j}$ and propagating the resulting changes to the model’s predictions.
However, CBMs are prone to a phenomenon known as information leakage (Margeloiu et al., 2021; Mahinpei et al., 2021), where the representation $\bm{z}$ encodes input information that cannot be attributed to the predefined concepts $\bm{c}$. We refer to this additional information as nuisances $\bm{n}$. Information leakage raises two main concerns: (i) it undermines interpretability, since $z_{j}$ cannot be fully explained by its corresponding concept $c_{j}$; and (ii) it compromises the validity of interventions—modifying $z_{j}$ may alter not only the associated concept $c_{j}$, but also other unintended information encoded in $z_{j}$.
We argue that information leakage stems from a fundamental limitation in the current formulation of CBMs: the absence of an explicit Information Bottleneck (IB) (Tishby et al., 2000) that actively constrains $z_{j}$ to exclude information unrelated to $c_{j}$. While the second objective in CBMs encourages each $z_{j}$ to retain $c_{j}$ in its entirety, it does not enforce that $z_{j}$ captures only information about $c_{j}$. In the worst-case scenario, $z_{j}$ could encode the entire input $\bm{x}$ and still satisfy this objective.
@@
To address this issue, we propose Minimal Concept Bottleneck Models (MCBMs), which incorporate an IB into each $z_{j}$. This ensures that $z_{j}$ not only retains all the information about its associated concept $c_{j}$, but also excludes any information unrelated to $c_{j}$, as summarized in Table 1. The name reflects that $z_{j}$ is trained to be a minimal sufficient statistic of $c_{j}$ (Fisher, 1922; 1935), in contrast to traditional CBMs where $z_{j}$ is optimized to be merely a sufficient statistic of $c_{j}$. As illustrated in Figure 1, this design yields disentangled representations that directly address the two shortcomings of CBMs: (i) it improves interpretability, since $z_{j}$ can be fully explained by its corresponding concept $c_{j}$; and (ii) it enables valid interventions—modifying $z_{j}$ affects only the associated concept $c_{j}$.
@@ 2.1 Data Generative Process
For this scenario, we consider inputs $\bm{x}\in\mathcal{X}$, targets $\bm{y}\in\mathcal{Y}$, concepts $\bm{c}=\{c_{j}\}_{j=1}^{m}\in\mathcal{C}$ and nuisances $\bm{n}\in\mathcal{N}$, such that $p(\bm{x},\bm{y},\bm{c},\bm{n})=p(\bm{x}|\bm{c},\bm{n})p(\bm{y}|\bm{x})p(\bm{c},\bm{n})$, i.e., the inputs $\bm{x}$ are described by the concepts $\bm{c}$ and the nuisances $\bm{n}$, and the targets $\bm{y}$ are fully described by the input $\bm{x}$. The only difference between $\bm{c}$ and $\bm{n}$ is that the former are observed, while the latter are not or, in other words, labels on $\bm{c}$ are provided to us. We assume access to a training set $\{\bm{x}^{(i)},\bm{y}^{(i)},\bm{c}^{(i)}\}_{i=1}^{N}$, which defines the empirical distribution $p(\bm{x},\bm{y},\bm{c})=\sum_{i=1}^{N}\delta\left(\bm{x}-\bm{x}^{(i)}\right)\delta\left(\bm{y}-\bm{y}^{(i)}\right)\delta\left(\bm{c}-\bm{c}^{(i)}\right)$. The graphical model corresponding to this generative process is shown in Figure 2(a) for the case of two concepts.
@@ 2.2 Vanilla Models
In machine learning, the most commonly studied problem is that of predicting $\bm{y}$ from $\bm{x}$, which serves as a foundation for more specialized tasks. We refer to models trained to address this problem as Vanilla Models (VMs). These models typically operate by first extracting an intermediate representation $\bm{z}\in\mathcal{Z}$ from the input $\bm{x}$ via an encoder $p_{\theta}(\bm{z}\mid\bm{x})$ parameterized by $\theta$. Subsequently, a prediction $\hat{y}\in\mathcal{Y}$ is produced from $\bm{z}$ using a task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$ parameterized by $\phi$. Since $\bm{z}$ is intended to facilitate accurate prediction of $\bm{y}$, the mutual information between $\bm{z}$ and $\bm{y}$, denoted $I(Z;Y)$, should be maximized. In Appendix B.1, we formally show that:
@@
Figure 2(b) shows the graphical model of a Vanilla Model with a two-dimensional representation $\bm{z}$. Black edges represent the encoder $p_{\theta}(\bm{z}\mid\bm{x})$, while green edges indicate the task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$. The encoder is typically chosen to be deterministic, i.e., $p_{\theta}(\bm{z}\mid\bm{x})=\delta\left(\bm{z}-f_{\theta}(\bm{x})\right)$, where $f_{\theta}:\mathcal{X}\to\mathcal{Z}$ is a neural-network–parameterized mapping. However, for tractability reasons (see Section 2.4), we adopt a stochastic formulation where $p_{\theta}(\bm{z}\mid\bm{x})=\mathcal{N}\left(\bm{z};f_{\theta}(\bm{x}),\sigma_{x}^{2}I\right)$, as summarized in Table 2. The choice of task head $q_{\phi}(\hat{\bm{y}}\mid\bm{z})$ depends on the structure of the output space $\mathcal{Y}$, also detailed in Table 2. The objective in Equation 1 corresponds to minimizing the cross-entropy loss when $\bm{y}$ is binary or multiclass, and to minimizing the mean squared error between $\bm{y}$ and $g^{y}_{\phi}(z)$ when $\bm{y}$ is continuous.
@@ 2.3 Concept Bottleneck Models
    Concepts $c_{j}\in\bm{c}$, are assumed to be binary. If a concept is originally multiclass with $k$ categories, it is converted into $k$ binary concepts (One-vs-Rest (Rifkin and Klautau, 2004)).
@@
Since $\sigma$ is invertible, this setup permits defining $p(z_{j}\mid c_{j})\approx\sigma^{-1}(c_{j})$. However, this is ill-defined at the binary extremes, as $\sigma^{-1}(1)=-\sigma^{-1}(0)=\infty$. To address this, in practice, $p(z_{j}\mid c_{j}=0)$ and $p(z_{j}\mid c_{j}=1)$ are set as the 5th and 95th percentiles of the empirical distribution of $z_{j}$, respectively (Koh et al., 2020). This workaround, however, introduces two crucial issues discussed in Section 5.
Table 2: Distributions considered in this work. $f_{\theta}:\mathcal{X}\to\mathcal{Z}$ is typically a large neural network while $g^{y}_{\phi}:\mathcal{Z}\to\mathcal{Y}$, $g^{c}_{\phi}:\mathcal{Z}\to\mathcal{C}$ and $g^{z}_{\phi}:\mathcal{C}\to\mathcal{Z}$ are comparatively lightweight networks (see Appendix F.1). Throughout this work, we model $\bm{z}$ as a continuous latent representation.
@@ 2.4 Minimal Concept Bottleneck Models
    For categorical concepts, $g^{z}_{\phi}(c_{j})=\lambda\cdot\mathrm{one\_hot}(c_{j})$. This mirrors the Prototypical Learning (Snell et al., 2017) approach where class-dependent prototypes $\{g^{z}_{\phi}(c_{j})\}_{j=1}^{m}$ are fixed.
@@
To optimize MCBMs, we combine the three previously introduced objectives, incorporating two key considerations. First, we approximate the expectations over $p(\bm{x},\bm{y})$ and $p(\bm{x},c_{j})$ using the empirical data distribution, replacing integrals with summations over the dataset. Second, to enable gradient-based optimization through the stochastic encoder, we apply the reparameterization trick (Kingma, 2013): $\mathbb{E}_{p_{\theta}(\bm{z}|\bm{x})}\left[\log q_{\phi}(\bm{y}|\bm{z})\right]\approx\sum_{i}\log q_{\phi}\left(\bm{y}|f^{\prime}_{\theta}\left(x,\epsilon^{(i)}\right)\right)$ and $E_{p_{\theta}(z_{j}|\bm{x})}\left[\log q_{\phi}(\hat{c}_{j}|z_{j})\right]\approx\sum_{i}\log q_{\phi}\left(\hat{c}_{j}\big|f^{\prime}_{\theta,j}\left(\bm{x},\epsilon^{(i)}\right)\right)$, where $f^{\prime}_{\theta}(\bm{x},\epsilon)=f_{\theta}(\bm{x})+\sigma^{2}_{x}I\epsilon$ (due to the choice of $p_{\theta}(\bm{z}|\bm{x})$ in Table 2), $\epsilon\sim\mathcal{N}\left(0,I\right)$ and $f^{\prime}_{\theta,j}\left(x,\epsilon\right)$ corresponds to the element $j$ of $f^{\prime}_{\theta}\left(\bm{x},\epsilon\right)$. Combining the considerations above yields the final objective for training MCBMs, as shown in Equation 6, where $\beta$ and $\gamma$ are hyperparameters. The first term corresponds to the objective used in Vanilla Models, the second term is introduced in CBMs, and the third is specific to MCBMs. Detailed training algorithms for the various cases listed in Table 2 are provided in Appendix C.
@@
Figure 2: Graphical models of the different systems described for two concepts and two-dimensional representations. Appendix A shows the analogous figure for $m$ concepts and $m$-dimensional representations. Inputs $\bm{x}$ are defined by some concepts $\{c_{j}\}_{j=1}^{m}$ and nuisances $\bm{n}$; and targets $\bm{y}$ are defined by $\bm{x}$ (gray arrows). Vanilla models obtain the representations $\{z_{j}\}_{j=1}^{m}$ from $\bm{x}$ through the encoder $\color[rgb]{0,1,1}p_{\theta}(\bm{z}|\bm{x})$ (cyan arrows) and solve the task $\hat{\bm{y}}$ sequentially through the task head $\color[rgb]{0,1,0}q_{\phi}(\hat{\bm{y}}|\bm{z})$ (green arrows). Concept Bottleneck Models make a prediction $\hat{c}_{j}$ of each concept $c_{j}$ from one representation $z_{j}$ through the concept head $\color[rgb]{0,0.49609375,1}q_{\phi}(\hat{c}_{j}|z_{j})$ (blue arrows). Minimal CBMs make a prediction $\hat{z}_{j}$ of each representation $z_{j}$ from one concept $c_{j}$ through the representation head $\color[rgb]{1,0,0}q_{\phi}(\hat{z}_{j}|c_{j})$ (red arrows).
@@ 3 Related Work
Information leakage occurs when the learned representation $\bm{z}$ encodes information outside the concept set $\bm{c}$, reducing both interpretability and intervenability (Margeloiu et al., 2021; Mahinpei et al., 2021). This arises when the Markovian assumption fails—i.e., when the target $\bm{y}$ is not fully determined by the concept set $\bm{c}$, or equivalently, $p(\bm{y}|\bm{c})\neq p(\bm{y}|\bm{c},\bm{x})$, which is typically the case in real-world scenarios (Havasi et al., 2022). Concept Embedding Models (CEMs) (Espinosa Zarlenga et al., 2022) were proposed to mitigate the accuracy–interpretability trade-off in CBMs. However, this trade-off is fundamentally limited by the chosen concept set, and CEMs may even be less interpretable than standard CBMs, as their more entropic representations tend to amplify information leakage, a critique formalized in (Parisini et al., 2025). Havasi et al. (2022) also introduced Hard Concept Bottleneck Models (HCBMs), which predict $\bm{y}$ from binarized concept predictions $\hat{c}_{j}$ rather than from $\bm{z}$ (see Appendix D), thereby imposing an ad-hoc Information Bottleneck. Extensions such as Autoregressive CBMs (ARCBMs) and Stochastic CBMs (SCBMs)(Havasi et al., 2022; Vandenhirtz et al., 2024) incorporate dependencies between concepts. Energy-based CBMs (Xu et al., 2024a) replace the task head with an energy function that scores concept–label compatibility, enabling richer and more structured concept relationships. Beyond architectural changes, prior work has characterized leakage using information-theoretic quantities (Parisini et al., 2025; Makonnen et al., 2025). To the best of our knowledge, this is the first work that leverages Information Theory to (i) formally identify the underlying design flaw—CBMs require each $z_{j}$ to predict $c_{j}$ but do not restrict $z_{j}$ from encoding additional nuisance information—and (ii) introduce a principled, variational IB objective that directly constrains each latent variable to retain only concept-relevant information.
@@
The Information Bottleneck (IB) (Tishby et al., 2000) provides a principled way to balance preserving information about a factor with compressing the representation. In this framework, a representation is sufficient if it retains all the information about the variable of interest, and minimal if it contains only that information (Achille and Soatto, 2018a; b; Shwartz Ziv and LeCun, 2024). Computation of these quantities is intractable, so variational methods have been developed to derive tractable evidence bounds (Alemi et al., 2016; Fischer, 2020).